OpenStudy (aonz):

Find the tangent to each curve at the point t=3 for x=5t^2 y=10t

OpenStudy (aonz):

i got 3 as gradient sorry

OpenStudy (aonz):

ill show my working

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{\frac{dy}{dt}}{\frac{dx}{dt}} \]

OpenStudy (aonz):